# Filamentary Hierarchies and Superbubbles I: Characterizing filament properties across a simulated spiral galaxy

### v2

@@ 1 Introduction
In the present paper, we build upon and mine the data in the Zhao et al. (2024) simulations to characterize, at a single snap-shot in time, filament properties on the galactic scales across the entire simulated spiral galaxy. Our goal is to provide a complete, statistically significant analysis of various filament properties on scales $>$20 pc including probability distribution functions of their masses, lengths, line masses and other related physical properties such as the relative orientation of magnetic field lines. The associated velocity fields and dynamics of filamentary and accretion flows in relation to forming clusters and magnetic field dynamics will be investigated in two follow-up papers in this series (M. Wells et al. 2025, submitted; R. Pillsworth et al. 2025, in prep). Our analysis features a method of filament extraction using the FilFinder code (Koch & Rosolowsky, 2015) that has been used successfully in observational studies of isolated, smaller scale filaments (see Gu et al., 2024; Williams et al., 2024; Zhang et al., 2024; Zhou & Davis, 2024, for some recent works). This method is robust across scales, and allows a standardized treatment of filamentary structure for both observations and simulations.
@@ 3 Data & Methods
The numerical data that we use in this paper come from the simulations of a Milky Way type galaxy from Zhao et al. (2024), which are run in ramses with the AGORA project initial conditions (Kim et al., 2016). These include a dark matter halo with MDM halo = 1.074 x $10^{12}$ Mdirect-$\odot$, RDM halo = 205.5 kpc, and a circular velocity of vc,DM halo = 150 km/s, an exponential disk with Mdisk = 4.297 x $10^{10}$ Mdirect-$\odot$, and a stellar bulge with Mbulge = 4.297 x $10^{10}$ Mdirect-$\odot$ (Kim et al. (2016)). For full details of the simulation setup we refer the reader to Zhao et al. (2024), which outlines all the details of the simulation.
@@ 4.1 Filament Width Analysis
Overall, while the Gaussian fitting of FilFinder is largely successful and results in few of the structures being culled, the same is not true for RadFil. We find that, due to the complex morphologies of many of these galactic filamentary structures (see Figure 1 for examples), fitting widths using RadFil resulted in $<$5% of our total filament population surviving. This is likely because RadFil is designed to perform on a filament-by-filament basis, instead of across a large population of filaments. This would have necessitated fine tuning the width profiles of each of our 500 total filaments. Comparing our few RadFil profiles to the Gaussian profiles fit by FilFinder, we found no difference in the measured widths when accounting for errors. Given the small surviving population, we decided to perform the rest of our analysis based only on the Gaussian widths found by FilFinder. In future work analyzing filamentary structures in specific highly resolved regions discussed in Zhao et al. (2024), we will implement RadFil’s profile fitting on a more individual basis to directly compare Gaussian vs. Plummer profiles for filament widths. This future work will also draw direct comparison to the limited Plummer profile fits presented in Zhao et al. (2024).
@@ 4.2 Filament Population Statistics
In Figures 4 and 5, we plot the length and mass distributions, respectively, of all of the filaments. We show distributions for our culled sample, the filaments that pass the cuts described in the previous section (length $>$25 pc, robust width measurements) – totaling 325 filaments. We calculate a total mass using the width of the filament as identified by the Gaussian fit.
@@
Figure 5: Histogram of the filament masses with our fitted power-law model across the entire mass range – $5\times 10^{4}$ $5\times 10^{8}$ M shown with the black line. Green dotted and dot-dashed lines should the median and mean mass of the filaments, respectively.
@@ 4.3 Theoretical and Measured Line Masses
In Figure 6 we provide both the measured and critical average line masses of our filament sample. From left to right, we plot measured line masses against increasingly detailed theoretical line masses, going from the simple thermal sound speed contribution (Equation 1) calculated using the local sound speed along the filament, to the full velocity dispersion (including non-thermal motions) in the middle panel (Equation 3), to the magnetically corrected total velocity dispersion – calculated using local velocity dispersion and Alfven speeds – in the rightmost panel. In each plot, we also include a line along a line mass ratio of 1, representing the transition value from subcritical ($<$ 1) to supercritical ($>$ 1) filaments.
@@ 5.1 Filaments in context
In Figure 8, we plot some key properties of our simulated filaments in the context of the plethora of observational filament data compiled in Hacar et al. (2023, their figure 2). The publicly available subset of Hacar et al. (2023) data points are shown as grey stars, including the data from Schisano et al. (2020). For additional context, especially for large scale filaments, we also include the original, extended and highly extended measurements of the Nessie filament, discussed in Jackson et al. (2010) and Goodman et al. (2014), as well as both the dense and total gas mass of the Maggie filament (Syed et al., 2022). For our own data, we distinguish the average subcritical (orange x’s) from the supercritical (purple diamonds) filaments, where we use the ‘total$+$magnetic’ critical line mass to determine filament criticality.
@@ 5.2 Mechanisms of Large-Scale Filamentary Structure Formation
To demonstrate that these are indeed coherent structures, we show in Appendix B the 3D spines of these filaments in Figures 15-18. These 3D spines are extrapolated from the 2D spines found by FilFinder using the cell of maximum density along z at each (x,y) coordinate of the spine. We then measure the velocity gradient along these structures - arguing that velocity coherence is a good measure of the structure as a filament. All four filaments we use as examples are continuous structures in 3D and exhibit smooth velocity gradients with values consistent with the cold neutral medium, which we discuss in more detail in Appendix B.
@@ 5.4 Magnetic Field Orientations
We plot the distribution of the average values of $|\mathrm{cos}(\theta)|$ between the magnetic field and filaments in Figure 13. Soler et al. (2013) find a positive correlation between the angle between the filament and the magnetic field and column density, and thus we plot the distribution for all of our filaments as well as for the 50th, 75th and 95th percentile of average filament column density. We also include a vertical dashed line meant to demarcate parallel orientations; i.e., as those where $|\mathrm{cos}(\theta)|$geq 0.5.
@@ Appendix B Velocity Coherence of Large scale Filaments
We verify the structure of our four large-scale filaments highlighted in Figure 1 and §5.2 by analysis of the velocity structure of them. The analysis is performed in two primary steps: checking the velocity contours around the structures and verifying the spines in 3D position space. We show the results of these analysis steps in Figures 15-18.
The 3D spines are constructed by employing a simple peak-finding method in density along the z-axis, limited to the range of the disk of the Galaxy. The spines are ordered in 3D using a B-spline interpolation method from SciPy (Virtanen et al., 2020). Importantly, Figures 15-18 showcase a still image of the 3D spines of these structures that has been smoothed for better visualization. However, the connectedness of the structures along z is first verified with no smoothing. The smoothing parameter is defined as the maximum allowed weighted difference between the spline point and the actual coordinate of the line where, in our case, the weights are the coordinates in cm along each dimension. For a smoothing parameter, s, set to up to 30000, we find no significant difference in the structures’ connectivity. We also do not interpolate between points, maintaining the same number of points along each structure to further mitigate any ‘false’ sub-structure which could appear due to the B-spline method.
Each of the four filaments presented in Figures 15-18 indeed have a persistent 3D structure. The velocity gradient along the 3D spines also confirms the velocity coherence of these structures, showing very few changes in the gradient along the spines. The few places with higher ($\leq$12 km/s/pc) velocity gradients either remain within reasonable values of the cold ($\sim$50K), mostly atomic gas of these filaments, or appear at areas of significantly higher density in which we expect massive clusters are beginning to form (see, for instance, Zhao et al. (2024) for a discussion of the clumps in their ‘active’ region, which is contained in our IS filament of Figure 15).
@@ Appendix C Column density contours
We verify the global column density cutoff given to the masking step of FilFinder by visualizing the final chosen column density threshold (0.003 g cm-2 = 3×$\times 10^{21}$ cm-2) on the extreme ends of our mass distribution. For both the lowest and highest mass filaments in our final dataset, the structures are well contained by the chosen column density cutoff. This further confirms that we accurately probe structures in the cold, neutral medium and are not tracing much more diffuse, atomic structures. We do note that the mask does not limit the width fitting or branch algorithms of FilFinder, such that filament widths or branches from main filaments can extend past the edges of the mask and, thus, sample some lower density data. However, these facts do not largely affect our analysis. We do not analyze any branches, limiting all of our analysis only to the main filaments identified in the skeletonization step. Additionally, in our data the vast majority of filament widths are within the bounds of the mask applied by FilFinder. Any filaments that may extend past the edges of the mask are those which are underresolved anyways, and are culled from our sample due to high errors on the width measurement.

### v3

@@ 1 Introduction
In the present paper, we build upon and mine the data in the Zhao et al. (2024) simulations to characterize, at a single snap-shot in time, filament properties on the galactic scales across the entire simulated spiral galaxy. Our goal is to provide a complete, statistically significant analysis of various filament properties on scales >>20 pc including probability distribution functions of their masses, lengths, line masses and other related physical properties such as the relative orientation of magnetic field lines. The associated velocity fields and dynamics of filamentary and accretion flows in relation to forming clusters and magnetic field dynamics will be investigated in two follow-up papers in this series (M. Wells et al. 2025, submitted; R. Pillsworth et al. 2025, in prep). Our analysis features a method of filament extraction using the FilFinder code (Koch & Rosolowsky, 2015) that has been used successfully in observational studies of isolated, smaller scale filaments (see Gu et al., 2024; Williams et al., 2024; Zhang et al., 2024; Zhou & Davis, 2024, for some recent works). This method is robust across scales, and allows a standardized treatment of filamentary structure for both observations and simulations.
@@ 3 Data & Methods
[…]
@@ 4.1 Filament Width Analysis
Overall, while the Gaussian fitting of FilFinder is largely successful and results in few of the structures being culled, the same is not true for RadFil. We find that, due to the complex morphologies of many of these galactic filamentary structures (see Figure 1 for examples), fitting widths using RadFil resulted in <<5% of our total filament population surviving. This is likely because RadFil is designed to perform on a filament-by-filament basis, instead of across a large population of filaments. This would have necessitated fine tuning the width profiles of each of our 500 total filaments. Comparing our few RadFil profiles to the Gaussian profiles fit by FilFinder, we found no difference in the measured widths when accounting for errors. Given the small surviving population, we decided to perform the rest of our analysis based only on the Gaussian widths found by FilFinder. In future work analyzing filamentary structures in specific highly resolved regions discussed in Zhao et al. (2024), we will implement RadFil’s profile fitting on a more individual basis to directly compare Gaussian vs. Plummer profiles for filament widths. This future work will also draw direct comparison to the limited Plummer profile fits presented in Zhao et al. (2024).
@@ 4.2 Filament Population Statistics
In Figures 4 and 5, we plot the length and mass distributions, respectively, of all of the filaments. We show distributions for our culled sample, the filaments that pass the cuts described in the previous section (length >>25 pc, robust width measurements) – totaling 325 filaments. We calculate a total mass using the width of the filament as identified by the Gaussian fit.
@@
Figure 5: Histogram of the filament masses with our fitted power-law model across the entire mass range – $5\times 10^{4}$ - $5\times 10^{8}$ M shown with the black line. Green dotted and dot-dashed lines should the median and mean mass of the filaments, respectively.
@@ 4.3 Theoretical and Measured Line Masses
In Figure 6 we provide both the measured and critical average line masses of our filament sample. From left to right, we plot measured line masses against increasingly detailed theoretical line masses, going from the simple thermal sound speed contribution (Equation 1) calculated using the local sound speed along the filament, to the full velocity dispersion (including non-thermal motions) in the middle panel (Equation 3), to the magnetically corrected total velocity dispersion – calculated using local velocity dispersion and Alfven speeds – in the rightmost panel. In each plot, we also include a line along a line mass ratio of 1, representing the transition value from subcritical (<< 1) to supercritical (>> 1) filaments.
@@ 5.1 Filaments in context
In Figure 16, we plot some key properties of our simulated filaments in the context of the plethora of observational filament data compiled in Hacar et al. (2023, their figure 2). The publicly available subset of Hacar et al. (2023) data points are shown as grey stars, including the data from Schisano et al. (2020). For additional context, especially for large scale filaments, we also include the original, extended and highly extended measurements of the Nessie filament, discussed in Jackson et al. (2010) and Goodman et al. (2014), as well as both the dense and total gas mass of the Maggie filament (Syed et al., 2022). For our own data, we distinguish the average subcritical (orange x’s) from the supercritical (purple diamonds) filaments, where we use the ‘total++magnetic’ critical line mass to determine filament criticality.
@@ 5.2 Mechanisms of Large-Scale Filamentary Structure Formation
To demonstrate that these are indeed coherent structures, we show in Appendix B the 3D spines of these filaments in Figures 18-21. These 3D spines are extrapolated from the 2D spines found by FilFinder using the cell of maximum density along z at each (x,y) coordinate of the spine. We then measure the velocity gradient along these structures - arguing that velocity coherence is a good measure of the structure as a filament. All four filaments we use as examples are continuous structures in 3D and exhibit smooth velocity gradients with values consistent with the cold neutral medium, which we discuss in more detail in Appendix B.
@@ 5.4 Magnetic Field Orientations
We plot the distribution of the average values of $|\mathrm{cos}(\theta)|$ between the magnetic field and filaments in Figure 13. Soler et al. (2013) find a positive correlation between the angle between the filament and the magnetic field and column density, and thus we plot the distribution for all of our filaments as well as for the 50th, 75th and 95th percentile of average filament column density. We also include a vertical dashed line meant to demarcate parallel orientations; i.e., as those where $|\mathrm{cos}(\theta)|$ $\geq$ 0.5.
@@ 6 Conclusions
Corrigendum
In this corrigendum, we address some corrections to our recently published paper,Pillsworth et al. (2025) entitled “Filamentary Hierarchies and Superbubbles. I. Characterizing Filament Properties across a Simulated Spiral Galaxy”. In particular, due to an undetected error in the treatment of filament cutouts, our filament mass calculations need correction. The corrections affect Figures 5, 6 and 8 from Pillsworth et al. (2025) but we find no change to our main results. Additionally, we find no other analysis is affected by these corrections and the characteristics of the individual filaments we highlight in the original paper remain the same with these corrections applied to our code.
In detail, a small bug in the FilFinder code mainly affected the masses of filaments around $10^{6}$ M and lower. In saving cutouts (stamps) around individual filaments, the defaults in FilFinder accounted for a 20 pixel padding around the ends of the filaments in order to avoid accidentally cutting out sections of the filamentary structure. However, in testing an update to the FilFinder code, we found that this padding affected the coordinates returned for the beginning of the filament in these cutouts. As such, our calculations of mass were using a skewed or misplaced filament spine that did not quite coincide with the actual filament identified.
We found that this bug caused a significant change in the measured masses of shorter (and therefore lower mass) filaments because of two main properties:
  1. 1.
    a minimum 20 pixel offset represents a larger fractional percentage of shorter filaments than our kpc-scale filaments,
  2. 2.
    the shorter filaments tend to be closer to superbubble structures, which have dense shock fronts expanding outwards but are not classified as filaments.
The combination of these issues resulted in our masses being overestimated for shorter filaments—sometimes by as much as 3 orders of magnitude— so that short filaments that originally had masses of approximately $10^{6}$ in our published paper are now reduced to the much more reasonable $10^{3}$ M range. We have patched the code by setting the default pad to 0 pixels, effectively skipping over the padding. We find that for our data with a working resolution of 5.2 pc, the lack of padding does not result in any structure being cut off from our filament stamps.
Furthermore, we verified this update in the calculation of the masses by selecting a handful of filaments at our low-mass end and calculating their masses by hand using the projected cell masses from our simulation data. We find these calculated masses to be consistent with those calculated in our updated analysis. With the correction for this issued now verified, the fix will be included in the next release of FilFinder.
In the top panel of Figure 14 we present our revised mass distribution of the filaments in our sample, replacing Figure 5 of Pillsworth et al. (2025). We note that filament masses now descend to 1000 M, and this results in a much better fit across the entire range of filament masses - changing the power-law relation from the previously reported 1.85 to 1.24.
In the bottom panel of Figure 14, we replot our revised data and include in red the PDF of the observed Milky Way molecular clouds identified by Rice et al. (2016). We find the latter has a power-law in our mass PDF formulation of 1.47, slightly greater than our revised filament PDF with a value of 1.24. This updated power-law index for the data reported in Rice et al. (2016) arises because of the differences in the definitions of the PDFs between the two works.
Figure 14: Top: Our filament mass distribution with corrected masses. The dotted and dash-dotted lines represent the median and mean mass of filaments, respectively. We show our new power-law fit in black, corresponding to a power-law index of 1.24. Bottom: The same mass histogram, as well as the Milky Way cloud masses (in red) from Rice et al. (2016). The addition of the dashed black line shows the power-law fit to the cloud catalogue, with a power-law index of 1.47 as measured by us.
These corrections to our mass calculations also affect the general line-mass trends in our filament population. We show joint plots of the theoretical and measured line masses of our filament population in Figure 15 which replaces Figure 6 of Pillsworth et al. (2025). Moving from the left to the right panels in Figure 15, we provide an increasingly detailed calculation of theoretical line mass, starting with the thermal line mass defined only by the sound speed (left), then including the total non-thermal velocity dispersion (center) and finally including the Alfven speed to the theoretical calculation (right). Each panel also shows the line of equality, delineating the subcritical and supercritical populations.
The largest change we note is the significant shift to subcritical line-mass ratios when the magnetic correction to the theoretical line mass is included. Now 15% of our filaments are supercritical when we consider the effects of the Alfven speed on the stability of filaments. We note no significant changes in the line-mass ratios of the thermal and velocity dispersion correction to the line-masses. The fact that magnetic support of filaments makes them more subcritical may align with star formation efficiencies expected in a Milky Way-like galaxy simulation, which is approximately 10% in our data. However, more analysis is needed to determine the relation between filament supercriticality and star formation in a galactic disk.
Figure 15: Theoretical vs. measured line mass joint plots for the filaments in our population, with corrected mass calculations. From left to right, we show the thermal critical line mass, the total critical line mass including velocity dispersion and the total critical line mass with a magnetic field correction. The black line in each shows the critical ratio of 1, separating the plane into sub-critical and supercritical regions.
Finally, we update the results of Figure 8 in Pillsworth et al. (2025) in Figure 16 below. We show a compilation of observed filaments from public data of Hacar et al. (2023), including that of Schisano et al. (2020), as well as our subcritical (yellow X’s) and supercritical (purple diamonds) filaments in a line-mass plot. In black X’s we show three mass and length measurements of the Nessie (G339) filament from Jackson et al. (2010); Goodman et al. (2014). Magenta X’s show the measurements of the dense and the dense+diffuse gas mass of the Maggie filament (Syed et al., 2022). In addition, we simplify the line-mass relations we include in comparison to the original. The solid and dashed pale red lines show the thermal line-mass scaling for sound speeds of 0.2 and 2.0 km/s, respectively. We include the length-scaled velocity dispersion relation as derived in Hacar et al. (2023) in its original form in the black line, as well as with $L_{0}=5.0$ pc to match our physical resolution in the blue line. Finally the 5.0 pc scaled line with a magnetic correction is shown in green.
With the corrected mass range, the most notable change is the clear overlap between our simulation data and the collected observational data. We also note that our filament population now tends to match the length-scaled relation fit by Hacar et al. (2023), with an adjustment to the scale length from 0.5 to 5.0 pc, in line with our resolution of 5.2 pc in our data. This fit with the adjusted scale length continues to fit the simulated population well with the addition of a magnetic field correction, though the low-mass end in our filament population does not extend far enough to distinguish whether the magnetic correction becomes crucial to fit the trend correctly below $10^{4}$ M.
Importantly, our finding that there is no relation that can accurately separate subcritical and supercritical filaments remains valid. Even with our corrected masses, the only line that comes close to splitting the supercritical and subcritical relation is the thermal line scaling corresponding to a sound speed of 2.0 km/s, in line with sound speeds expected for the cold, neutral medium. Even still, we see a significant population of subcritical filaments to the right of this line, nominally in the supercritical region of the plot. Given the lengths and masses of these subcritical filaments, we emphasize our original result in Pillsworth et al. (2025) that the local variations of line mass along such large-scale filaments are crucial to understanding their fragmentation.
Figure 16: Filament line mass: grey stars represent public data from Hacar et al. (2023); Schisano et al. (2020), reproduced with permission from corresponding PIs. The yellow X’s and purple diamonds represent our sub- and supercritical filaments, respectively. We include the classical, extended and optimistic measurements of the Nessie (G339) filament from Goodman et al. (2014) in black X’s. The magenta X’s show the properties of the Maggie filament from Syed et al. (2022) for the dense ridge and the combined dense+diffuse gas mass measurement. Light red lines show the thermal critical line mass for a sound speeds of 0.2 km/s (solid) and 2.0 km/s (dashed). The solid black line represents the length-scaled relation fit in Hacar et al. (2023) for their measured $L_{0}$=0.5 pc and a sound speed of 0.2 km/s. In the blue, we show the same relation but with $L_{0}$= 5.0 pc to match our resolution. The green line represents the same parameters as the blue, with an added magnetic field correction with Alfven speed 0.4 km/s.
Overall, we highlight that the qualitative interpretation of our results does not change after including the corrections to our mass distribution. Large scale, kpc filaments are still capable of being, on average, in a subcritical state despite good evidence of fragmentation and star formation. The large population of subcritical filaments, especially at smaller scales reaching down to our 5 pc resolution limit, highlights deficiencies in estimating the fragmentation or star formation activity of a filament based on an average line-mass ratio alone. We find no clear relation exists in the line-mass diagram that cleanly separates subcritical and supercritical populations, indicating, perhaps unsurprisingly, that the criticality of a filament is more more dominated by local fluctuations which analytical theory based on infinite, uniform filaments, was never set up to describe.
That said, we note that our numerical results are still in good agreement with the general trend in the M-L relation arising from the the length-velocity dispersion relation derived in Hacar et al. (2023) and find these lines fit the trend of measured line masses of the filaments. Whether or not there is a finely tuned, critical line that cleanly separates average sub from super critical states mush await future observational studies - which would be highly valuable.
@@ Appendix B Velocity Coherence of Large scale Filaments
We verify the structure of our four large-scale filaments highlighted in Figure 1 and §5.2 by analysis of the velocity structure of them. The analysis is performed in two primary steps: checking the velocity contours around the structures and verifying the spines in 3D position space. We show the results of these analysis steps in Figures 18-21.
The 3D spines are constructed by employing a simple peak-finding method in density along the z-axis, limited to the range of the disk of the Galaxy. The spines are ordered in 3D using a B-spline interpolation method from SciPy (Virtanen et al., 2020). Importantly, Figures 18-21 showcase a still image of the 3D spines of these structures that has been smoothed for better visualization. However, the connectedness of the structures along z is first verified with no smoothing. The smoothing parameter is defined as the maximum allowed weighted difference between the spline point and the actual coordinate of the line where, in our case, the weights are the coordinates in cm along each dimension. For a smoothing parameter, s, set to up to 30000, we find no significant difference in the structures’ connectivity. We also do not interpolate between points, maintaining the same number of points along each structure to further mitigate any ‘false’ sub-structure which could appear due to the B-spline method.
Each of the four filaments presented in Figures 18-21 indeed have a persistent 3D structure. The velocity gradient along the 3D spines also confirms the velocity coherence of these structures, showing very few changes in the gradient along the spines. The few places with higher ($\leq$12 km/s/pc) velocity gradients either remain within reasonable values of the cold ($\sim$50K), mostly atomic gas of these filaments, or appear at areas of significantly higher density in which we expect massive clusters are beginning to form (see, for instance, Zhao et al. (2024) for a discussion of the clumps in their ‘active’ region, which is contained in our IS filament of Figure 18).
@@ Appendix C Column density contours
[…]